[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue May 16, 2017

Monica Bozeman - NOAA Federal monica.bozeman at noaa.gov
Tue May 16 17:21:44 UTC 2017

All, in an attempt to switch over the processing of the GOES-16 in AWIPS2
from scp to and LDM feed, sending of the GOES-16 data to AWIPS2 was turned
off in anticipation of the switchover. Unfortunately, some complications
have come up and this switch has still not taken place. The person who
oversees the code that processes the GOES-16 data into NAWIPS and AWIPS2 is
on AL today, and therefore you will not see GOES-16 in your AWIPS machines

Once the POC gets back, or the issues with the switch gets resolved with
NCO (whichever happens first), data should start flowing into your AWIPS
again. Sorry for the inconvenience.


Monica Bozeman
National Hurricane Center
monica.bozeman at noaa.gov

On Tue, May 16, 2017 at 3:48 PM, TSB Admin <nhc.tsbadmin at noaa.gov> wrote:

> Here is today's summary of NHC computer operations:
> --- Mark
> ----------
> *TAFB:*
> NCO made some configuration changes on the cumulus N-AWIPS workstation,
> but the workstation quickly froze again. Thanks to Nelsie for helping with
> the testing. NCO has some other things they are going to try next.
> More generally, there are ongoing issues with several of the TAFB N-AWIPS
> workstations since the upgrade to RHEL6. NCO is trying to address these but
> are short-staffed and progress remains slow. We expect the situation will
> improve when NCO fills the vacant position at NHC, probably in mid-June.
> Until then, continue to report problems as they occur.
> The plumber has been called for the west Men's restroom.
> *HSU:*
> *Other items:*
> *NHC RFCs:*
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170516/fe55e103/attachment.html 

More information about the Ncep.nhc.nco_contacts mailing list